OpenStudy (anonymous):

Quick question: Why is -3*-3 = 9, but -3^2=-9?

pooja195 (pooja195):

a negative times a negative is a postivie

OpenStudy (anonymous):

But isn't -3 to the second power still -3 times -3?

pooja195 (pooja195):

=−3^2 =−1*3^2 =−1*(3*3) =−9

OpenStudy (anonymous):

\[-3^2=-3\times 3=-9\] a la order o' operations \[(-3)^2=(-3)\times (-3)=9\]

pooja195 (pooja195):

o.O spanish?

OpenStudy (anonymous):

french, but a very common and obvious idiom in english

OpenStudy (anonymous):

\(-3^2=-(3^2)=-(3\cdot 3)=-9\) while \((-3)^2=(-3\cdot-3)=9\)

OpenStudy (anonymous):

I get it. Thanks, everyone.